OpenStudy (anonymous):

surveyor took the following measurements from two irregular shaped pieces of land. some of the lengths angles measurements are missing. round to the nearest minute. a) find the length of a. b) find the measurement of angle H. write answer as, for example, 68 (degrees), 9

OpenStudy (loser66):

cannot open

OpenStudy (anonymous):

do u get it @Loser66

OpenStudy (anonymous):

i think that is 50 because 80-30 is 50 right @Loser66

OpenStudy (anonymous):

yea im learning about that but i don't completely understand it @agent0smith

OpenStudy (anonymous):

@charlotte123

OpenStudy (charlotte123):

Oh dear - I did not study this concept yet >.>

OpenStudy (charlotte123):

@trillestmali Looks like @agent0smith Is getting to It! I am of no help on the Subject >.>

OpenStudy (anonymous):

can we use the pathyon theory to solve for a? @agent0smith

OpenStudy (anonymous):

okay thanks doe @charlotte123

OpenStudy (anonymous):

but isn't the triangle in the right upper corner a right triangle??

OpenStudy (charlotte123):

@trillestmali No need to thank me LOL Good Luck and thank you for understanding :)

OpenStudy (anonymous):

can we use sochatoa??

OpenStudy (ranga):

If you want to avoid the 25 degree & 45 minutes angle on the right corner avoid the right triangle altogether and answer both parts using the middle triangle. Answer part b) first as angle H = angle DGH (alternate interior angles). Then apply law of sines in the middle triangle to solve for a.

OpenStudy (ranga):

**25 degree & 54 minutes**

OpenStudy (anonymous):

okay

OpenStudy (agent0smith):

The arrows might help to see where the stuff in the law of sinces comes from  \[\Large \frac{ \sin G }{ 80 }= \frac{ \sin 49 }{ 62 }\]

OpenStudy (anonymous):

should i cross multiple for this

OpenStudy (agent0smith):

You can, but you should learn faster ways to solve problems. Just multiply both sides by 80.

OpenStudy (anonymous):

i got sin(G)(62)=60.38

OpenStudy (agent0smith):

Cross-multiplying isn't necessary. Skip a step and just multiply both sides by 80

OpenStudy (agent0smith):

\[\Large 80*\frac{ \sin G }{ 80 }= \frac{ \sin 49 }{ 62 }*80\] what happens to the 80/80 on the left?

OpenStudy (anonymous):

they cancel

OpenStudy (agent0smith):

Exactly. Learn from this... there's very rarely a need to cross multiply.

OpenStudy (agent0smith):

Equations are often solved faster w/o it

OpenStudy (anonymous):

so in the right side u get .012??

OpenStudy (agent0smith):

No, check you're in degrees.

OpenStudy (anonymous):

-.015?

OpenStudy (agent0smith):

No...

OpenStudy (anonymous):

no i got 56.85 is that right

OpenStudy (anonymous):

now*

OpenStudy (agent0smith):

Show me what you're doing. I have no idea.

OpenStudy (anonymous):

\[\sin 49*80 the 62 *80 then divide then \]

OpenStudy (agent0smith):

I can't follow that. Write it in normal words, like (sin49/62)*80 or something.

OpenStudy (agent0smith):

Make sure you know what you are doing... the stuff on the right.\[\Large \sin G = \frac{ \sin 49 }{ 62 }*80\]

OpenStudy (anonymous):

sine(49)(80)/(62)(80)

OpenStudy (agent0smith):

See above.

OpenStudy (anonymous):

i multiple 80 top and bottom

OpenStudy (agent0smith):

Why??

OpenStudy (anonymous):

i though what ever u do in the top u have to to it in the bottom

OpenStudy (agent0smith):

Take a careful look at this. \[\Large \frac{ \sin 49 }{ 62 }*80\]Do *exactly* what is shown.

OpenStudy (agent0smith):

That's when you're modifying a fraction, we are not doing that

OpenStudy (anonymous):

okay i don't know what im missing here im back to my old answer -1.23

OpenStudy (anonymous):

what did u get?

OpenStudy (agent0smith):

It doesn't matter what I get, if you can't find it yourself. Do this, and exactly this, on your calculator. Make sure you have your calculator in degrees. 80*(sin49)/62

OpenStudy (anonymous):

.97

OpenStudy (agent0smith):

Good.

OpenStudy (anonymous):

finally ahah

OpenStudy (agent0smith):

\[\Large \sin G = 0.97\]now find G.

OpenStudy (anonymous):

.02

OpenStudy (agent0smith):

How did you get that? did you find sin^-1(0.97)???

OpenStudy (anonymous):

no i did sin(.97)=.02

OpenStudy (agent0smith):

sinG=0.97 you have to take the inverse sin of 0.97 to find G.

OpenStudy (anonymous):

so it's 75.9

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

so now we can do 180-49-75.9 to get angle h

OpenStudy (agent0smith):

Yep then we gotta find a  \[\Large \frac{ a }{ \sin 55.1 }=\frac{ 62 }{ \sin 49 }\]